OpenStudy (anonymous):

find the limit of sequence ((sin^2 n) / 5) ^n

OpenStudy (anonymous):

Easier look of this problem : \[\ (\frac{ \sin ^{2} n }{ 5 }) ^ {n}\]

OpenStudy (kainui):

As it approaches what?

OpenStudy (anonymous):

infinity sorry I forgot that haha

OpenStudy (kainui):

Haha no problem. Well, first off, what's your guess and why? What do you know about the sine function?

OpenStudy (anonymous):

sine x approach infinity should have the value between -1 and 1 right ?

OpenStudy (anonymous):

and sine ^ 2 x should be between 0 and 1

OpenStudy (anonymous):

and now I'm stuck :D

OpenStudy (kainui):

Sure, so think about the denominator. And while you do that, go ahead and think about two specific cases of the sin^2(n) \[(\frac{ 0 }{ 5 })^n and(\frac{ 1 }{ 5 })^n\]

OpenStudy (anonymous):

hmm..... are they indeterminate form 0^inf and 1^inf ?? I'm not sure 0^inf is simply 0 though.

OpenStudy (kainui):

Well, not exactly. Since in this case I have chosen the numerator to be 0 or 1 and it is no longer dependent on the variable. So the first is really just 0 raised to a power, and it doesn't matter what power you raise 0 to, it's always 0. The second one is (1/5)^n, so it's actually not 1^inf at all. Keep thinking and guessing though, I like it.

OpenStudy (anonymous):

oh gawd it's not 1 ..... It's 1/5. So 1/5 raised to inf is .... zero ?

OpenStudy (kainui):

Yeah, so what about some other possible value sin^2(x) can have between 0 and 1? They're all fractions. And any number less than 1 raised to a power becomes a smaller number. So everything in the numerator is just oscillating between 0 and 1, but as we approach infinity, no matter what it is, it'll still be 1 or 0 really. So it's just like we're taking the limit as (1/5)^n approaches infinity, which is 0. It just sort of wobbles a little bit while we get there.

OpenStudy (kainui):

A graph is good: http://www.wolframalpha.com/input/?i=%28sin%5E2%28x%29%2F5%29%5Ex+from+0+to+2 in fact, I only picked the graph to show from 0 to 2 because it approaches 0 so quickly before getting really far at all. Wolfram's a great tool to help you learn too if you don't use it to just solve all your calculus problems and cheat.

OpenStudy (anonymous):

ok. got it all. Thank you very much !! :)

OpenStudy (anonymous):

One way to approach this is to write \[ \left | \frac {\sin^2(n) } 5 \right|^n\leq \frac 1 {5^n} \] and deduce.